OpenStudy (anonymous):

A sample of 100 cans of peas showed an average weight of 14 ounces with a standard deviation of 0.7 ounces. If the distribution is fairly normal, how many cans will weigh over 14.7 ounces? 16 8 none 32

OpenStudy (michele_laino):

the distance between 14.7 and 14, is: 14.7-14=0.7 namely the distance between 14.7 and 14 is one time the standard deviation. is it ok?

OpenStudy (michele_laino):

since 0.7 = 1*0.7=1* standard deviation

OpenStudy (michele_laino):

so: